OpenStudy (anonymous):

Twelve friends came to Guy’s birthday party. He received seven gifts and a total of $95 in cash for his birthday. He buys an airplane model kit for $9, and he wants to donate $8 to each of several local animal shelters. How much money will he have left if he donates to as many animal shelters as possible?

OpenStudy (anonymous):

95-9=86 9*9=81 86-81=5

OpenStudy (anonymous):

Total cash = $95 After buying the model kit = $95 - $9 = $86 To donate $8 to \(n\) animal shelters, he would have to split up his total into as many units of 8 as possible: \[\frac{86}{8}=10+\frac{6}{8}\] i.e. he has $6 dollars remaining after donating $8 to each shelter for 10 different shelters.

OpenStudy (anonymous):

siths you corrected my wrong answer haha my bad

OpenStudy (anonymous):

no worries :)